OpenStudy (leozap1):

A $6,000.00 principal earns 8% interest, compounded semiannually. After 35 years, what is the balance in the account ok so do I add 8% of the balance 35 times? (Once for each year)

terenzreignz (terenzreignz):

This is compound interest, and it's compounded semianually, it actually means the effective interest rate is 4% for every semi-annum, or half-year. So, if there are 35 years, how many half-years? :D

OpenStudy (leozap1):

70

terenzreignz (terenzreignz):

That's right :) Now, this is compound interest, so, the balance earns interest every time. What you said, which is, adding 8% of the balance every time, is technically what we do, but that sounds rather tedious. Here's a better way: The balance, or accumulated value, after n-years is given by \[\Large P\left(1+ \frac{r}{m}\right)^{mn}\] Where P is the present value, or principal r is the compounded rate of interest m is the number of divisions per year (in this case 2, since it's compounded semianually, or per half-year) n is the number of years. Just plug in, and you'll have your answer ^_^

terenzreignz (terenzreignz):

n is still 35 years okay? And don't forget that r should not be in percent, but in decimal: 8% = 0.08

OpenStudy (leozap1):

None of the answers are in that forum though

terenzreignz (terenzreignz):

P is 6000, so... we basically have \[\Large 6000(1.04)^{70}\]

OpenStudy (leozap1):

Were did you get 1.04?

terenzreignz (terenzreignz):

\[\Large 1+ \frac{r}m\] r = 0.08 since the rate is 8% m is 2, since it's compounded per half-year (semianually) So... \[\Large 1+ \frac{0.08}2= 1+0.04=1.04\]

OpenStudy (leozap1):

Oh

OpenStudy (leozap1):

Is there a easier way to multiply it out Here are the answers a. 22,800 b. 39,600 c. 88,712 d. 93,429

terenzreignz (terenzreignz):

No... I'm sure they won't make you do this if you didn't have a calculator at your disposal. And yes, the answer is in the choices...

OpenStudy (leozap1):

1.04^70 Do I multiply that out first?

terenzreignz (terenzreignz):

Sure. It doesn't really matter... but that does seem the most logical way.. (please tell me you have a calculator)

OpenStudy (leozap1):

Yes I do

terenzreignz (terenzreignz):

then... just key in the stuff, and you'll be set ^_^

OpenStudy (leozap1):

It will still take for ever

terenzreignz (terenzreignz):

Oh... not a scientific calculator? LOL no matter, just google 1.04^70

terenzreignz (terenzreignz):

And then multiply it to the principal, 6000

OpenStudy (leozap1):

I dotn have oen of them

terenzreignz (terenzreignz):

Yeah, google also does calculating.

terenzreignz (terenzreignz):

Just google 1.04^70

OpenStudy (leozap1):

I used online calculator and it say its only 15.571618 that cant be right

terenzreignz (terenzreignz):

Yes, of course, that's not it yet, you still have to multiply that to the principal, 6000, remember? :P

OpenStudy (leozap1):

Oh yah lol

OpenStudy (leozap1):

Its D. :)

terenzreignz (terenzreignz):

That's right ^_^ Good job...

OpenStudy (leozap1):

Thanks :)

terenzreignz (terenzreignz):

Might want to keep this in mind, in case of more questions like this: \[\Large P\left(1+ \frac{r}{m}\right)^{mn}\] But remember, a formula can get you only so far... and not very far if you don't know what it's for ^_^

OpenStudy (leozap1):

Can you help me with another one please A boat cost $92,000.00 and the depreciates in value by 15% per year. How much will the boat be worth after 10 years? a. 18,112.45 b. 78,200.00 c. 18,941.98 d. 69,000.00

terenzreignz (terenzreignz):

ohh... depreciate... it's similar, except the rate is in the negative instead of positive, since the value is decreasing. Still this formula : \[\Large P\left(1+ \frac{r}{m}\right)^{mn}\] But take r = -15% or -0.15 and m = 1 since it's yearly; ie; the year is not divided.

OpenStudy (leozap1):

I am still confused

OpenStudy (leozap1):

Sorry I just don't understnad

terenzreignz (terenzreignz):

Evaluate: P = 92000 r = -0.15 m = 1 n = 10 yrs

OpenStudy (leozap1):

Ok so then it would be 0.925^10

terenzreignz (terenzreignz):

m=1, okay? not 2. Because this is per year, not per half-year.

OpenStudy (leozap1):

Oh

OpenStudy (leozap1):

So 0.85 ^ 10?

terenzreignz (terenzreignz):

Yes, and when that's done, multiply that to the principal.

OpenStudy (leozap1):

= 0.19687440434072?

terenzreignz (terenzreignz):

and multiply to the original value.

OpenStudy (leozap1):

A?